<p>21/2/22 16:00</p>
<p>Call from main control room, beam is ON, probably has been for several hours judging by rate on xy total plot.</p>
<p>Changes made to online monitoring code by OH, include setting to only 1 DSSD, correct offsets from the pulser walkthrough, offline analysis capabilities and X/Y position correction. </p>
<p>Monitoring code restarted in CARME_SORT/build_no_output</p>
<p>17:31 Run 50 ended, run 51 started now beam is definitely in the ring. Tried to insert span zero schottky into aida2 correlation scalar, however rate was far too high.</p>
<p>18:03 Removed span zero schottky signal, started new run 52.</p>
<p>With beam ON we noticed some missing channels in the xy total plot. Aida 1 and 2 stat totals on spectrum browser show missing channels. Asic check on both to no effect. Fees power cycled.</p>
<p>19:57 Started run 53. Currently waiting for beam</p>
<p>21:20 Beam ON</p>
<p>1*L plots checked for aida 1, several channels showed no data (noise)</p>
<p>Slow comparator LEC threshold lowered to 0x32 to aida 1, checked 1*L plots and channels are now showing lots of noise. So those channels appear to be quiet using 0x64 threshold. (attachment 3)</p>
<p>Threshold increased back to 0x64</p>
<p>22:33</p>
<p>System checks OK -white rabbit fails on all. ~30 seperation</p>
<p>Temperatures OK</p>
<p>Statistics within range - new scalars added for aida02 and aida05</p>
<p>Data being written ~300 kb/sec</p>
<p>Leakage current ~3.6 uA</p>
<p> </p>
<p>00:39</p>
<p>System checks checks OK -white rabbit fails on all. ~42 seperation</p>
<p>Temperatures OK</p>
<p>Statistics within range with new scalars</p>
<p>Data being written 256 kb/sec</p>
<p>Leakage current ~3.615 uA</p>
<p>XY counts dominated by hot channels - most blue pixels are ~10 counts</p>
<p> </p> |